OpenStudy (anonymous):

a cheetah ran 300ft in 2.92 sec. what was the cheetahs average speed in miles per hour please explain

OpenStudy (anonymous):

this is why i prefer the metric system

OpenStudy (anonymous):

how is that helping me

OpenStudy (anonymous):

ok so we have 3600 seconds in an hour (60^2) and google told me there's 5280 foot in a mile

OpenStudy (anonymous):

ok im still confused

OpenStudy (anonymous):

well, if he ran 300ft in 2.92 seconds than he ran 300/2.92 ft/s so about 102.739726027 feet per second

OpenStudy (anonymous):

i got that 2

OpenStudy (anonymous):

i need miles per hour

OpenStudy (anonymous):

102 feet is 0.019 mile or so (102/5280). so he runs about 0.019 mile per second

OpenStudy (anonymous):

multiply that by 3600 and you get miles per hour

OpenStudy (anonymous):

\[1 ~yard=3 ft,1~mile=1760 ~yards=1760 *3=5280~ft\]

OpenStudy (anonymous):

oh i see what you did ok thanks

OpenStudy (anonymous):

\[average~speed=\frac{ 300 ft }{ 2.92 \sec }=\frac{ \frac{ 300 }{ 5280 } }{ 2.92 }miles/\sec=\frac{ 300 }{ 5280 *2.92 }*3600 m/hr=?\]

OpenStudy (anonymous):

i got it thanks guys